OpenStudy (anonymous):

What is an equation for a direct variation whose graph passes through the point (–5, 3)?

OpenStudy (anonymous):

i think it would be a fraction...?

OpenStudy (stamp):

Direct variation:\[\frac{x}{y}=\frac{-5}{3}\]Solve for y.

OpenStudy (anonymous):

x=-3/5y?

OpenStudy (anonymous):

@stamp

OpenStudy (stamp):

You solved for x. Try solving for y.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

y=-3/5x

OpenStudy (anonymous):

?

OpenStudy (stamp):

Test it. Consider point (-5, 3) \[y=-\frac{3x}{5}\]\[(3)=-\frac{3(-5)}{5}\]\[3=3\] It works. Your answer is correct, good job.

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

maybe this y=mx+b 3=m*5+b you can chose anything for m and b